OpenStudy (anonymous):

Mary is 14 years younger than her sister Jane. In four years, Jane will be 3 times older than Mary. How old is Jane now? 3 7 17 21

OpenStudy (anonymous):

consider x represents the age of mary and y represents the age of jane X= y - 14 y+4 = 3 (x+4) subtitute first eqn in the second y+4 = 3(y - 10) solving it y + 4 = 3y - 30 2y = 34 y = 17

OpenStudy (anonymous):

oh thanks